# Topologically protected negative entanglement

### v3

@@ Abstract
First, we showcase a model with two topological edge modes that intersect at an exceptional crossing. Unlike typical topological band crossings [76, 77, 78, 79, 80, 81, 82] where the topological modes just have to be energetically degenerate, here we require them to also coalesce i.e. become parallel. A candidate model is given by the following 4-band Hamiltonian [73]
@@
In Fig. 1, we present three distinct scenarios corresponding to different parameter combinations, focusing particularly on the overlap $\eta(k)$ between the middle two eigenstates which straddle the Fermi energy $E_{F}=0$ (dashed line). Open boundary conditions (OBCs) are taken only along the $y$ direction, such that $k$ remains a good quantum number. In Fig. 1(a) with intersecting Hermitian topological modes (blue), $\eta(k)=0$ due to the exact orthonormality of Hermitian eigenstates. In Fig. 1(b) which is non-Hermitian ($\delta\neq 0$), $\eta(k)$ remains essentially zero due to the substantial bulk gap. However, in the non-Hermitian case with topological modes [Fig. 1(c)], the topological edge modes (red) cross and coalesce, forming an exceptional point, as reflected by the saturated squared overlap of $\eta(0)=1$. Only for this exceptional topological case do we see negatively entanglement entropy $S_{A}$ [Fig. 1(d)]; for the previous two gapless and gapped cases of Figs.1(a,b), $S_{A}$ respectively grows/saturates with $L$ as expected from usual conformal field theory [83, 84, 85, 86, 87].
@@
$\rm{Re}[S_{A}]\sim-0.3399\log L\approx\left(\frac{1}{3}-\frac{2}{3}\right)\log L,$ (8)
@@
scaling in the exceptional topological case differs from the previously reported $S_{A}\sim-\frac{2}{3}\log L$ scaling for a linearly dispersive exceptional point [65, 66, 68]. This discrepancy is attributed to gapless non-exceptional gapless crossing [gray in Fig.1(c)], which contributes the usual $\frac{1}{3}\log L$ entanglement. As such, the negative entanglement from exceptional topological crossings can be easily overshadowed by other non-exceptional topological crossings, and is in this sense not necessarily robust  [^4] This is particularly so if the exceptional dispersion is square-root, as for some generalizations of our model  [73]..
Enhanced negative entanglement from topological flat bands.– We next showcase an alternative construction that results in far more robust negative entanglement. Departing from the exceptional topological crossing paradigm, the key idea is that the non-Hermitian skin effect (NHSE) localizes all states towards a common boundary, such that the states would exhibit extremely high overlap if they are furthermore macroscopically energetically degenerate, as in a topological flat band. From our results below, this holds true even if the flat bands are not strictly gapless or defective.
To realize topologically protected flat bands across an extended range of $k$, it suffices to consider a minimal 2-component 5552 topological bands are sufficient, since the NHSE pushes both of them onto the same boundary and makes them overlap significantly. Hamiltonian
with $k$-dependent asymmetric off-diagonal hoppings, where $a_{0},b_{0}$ and hopping distance $B>0$ are all real. Under $y$-direction OBCs, the NHSE results in boundary state localization with a skin depth of $-2/\log[(b_{0}-\cos k)^{B}/a_{0}]$ and dynamically induces a real energy spectrum from non-Bloch PT-symmetry [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101]. Almost-flat topological bands exist whenever
$|a_{0}(b_{0}-\cos k)^{B}|\leq t^{2},$ (10)
[…]
$\det[H_{\text{y-OBC}}(k)]=[a_{0}(b_{0}-\cos k)^{B}]^{L_{y}},$ (11)
implying that for $10<|a_{0}(b_{0}-\cos k)^{B}|<$ 1, the topological gap also exhibits a $\Delta\sim(\text{Const.})^{-BL_{y}}$ inverse exponential scaling to zero as $L_{y}\rightarrow\infty$ (as derived in Sect. II.A of the Supplement [73]), even though it never exactly closes. This implies exponentially dependence of band flatness with $BL_{y}$. Perfect gap closure is only possible for $|b_{0}|\leq 1$, and below we investigate these two cases separately.
Figure 2: Robustly negative entanglement entropy from the gapped topological flat bands of our 2-component Hamiltonian (Eq. (9)), for parameters $B=1,t=0.8,a_{0}=1$ (with $b_{0}=1.2\neq 1$ to open up the gap). (a) For small cylinder length $L_{y}=5$ and $y$-direction OBCs, the energy spectrum $E(k)$ of the topological edge states |$\ket{\psi_{m1}},\ket{\psi_{m2}}$ (bolded) exhibits a small but visible gap, but their overlap factor $\eta(k)$ already approaches unity. (b) Upon increasing $L_{y}$ to 25, a topological flat band with exponentially small gap is observed within the non-trivial regime prescribed by Eq.(10), with $\eta(k)\approx 1$ extremely closely. (c) With $y$-PBCs, the midgap flat band disappears and $\eta(k)$ deviates markedly from unity, even though the gap is still small. (d) For $y$-OBCs but not $y$-PBCs, the overlap $\eta(k)$ saturates very close to unity once $L_{y}\sim 10^{1}$. (e) The entanglement entropy scaling behavior $S_{A}$ for different $L_{y}$. Notably, as $L_{y}$ increases, $S_{A}$ decreases with $\log L$ more rapidly as $S_{A}\sim-(\alpha L_{y}+\beta)\log L$, with $\alpha\approx 0.6633$, $\beta\approx-4.1817$ according to obtained from numerical fitting (black). It also saturates at $S_{min}\sim-L_{y}$ when $L\gtrapprox L_{y}$.
1. Gapped topological flat bands ($b_{0}>1$). – Even though the topological bands are gapped, they become almost flat and touching as $L_{y}$ is increased, as depicted in Figs. 2(a,b). At large $L_{y}$, their gap becomes exponentially small within the topologically non-trivial region given by Eq. (10), where the overlap factor $\eta(k)\approx 1$. This suggests that states within an extensive continuum of $k$ closely approximate EPs, which is unexpected since the Hamiltonian $H(k,k_{y})$ in Eq. (9) does not inherently feature EP crossings. As a comparison, for periodic boundary conditions(PBCs) in the $y$ direction [Fig. 2(c)], the topological flat band is absent, and the overlap $\eta(k)$ does not approach $1$ even as $L_{y}$ increased to a large value of $25$ where the (bulk) band gap becomes quite narrow. For $y$-OBCs, even at very small $L_{y}\approx 5$ number of layers, $\eta(k)$ is already very close to one [Fig. 2(d)]; at larger $L_{y}$, $\eta(k)$ converges exponentially to 1 despite the system being physically gapped.
This strong topological flat band-induced non-orthogonality ($\eta\approx 1$) is manifested in a strongly negative bipartite entanglement entropy $S_{A}$. As shown in Fig. 2(e), $S_{A}$ scales negatively with the cylinder circumference $\log L$, with a gradient that grows with its length $L_{y}$. From Sect. II.C of the Supplement [73], the exact dependence is established as $S_{A}\sim-(\alpha L_{y}+\beta)\log L$, where $\alpha\approx 0.6633$, $\beta\approx-4.1817$ as obtained from numerical fitting. Notably, this linear dependence on $L_{y}$ does not arise trivially because the length of the entanglement cut scales with $L_{y}$, since it is contributed only by the topological modes whose number do not scale extensively with system length. Rather, it arises because the band flatness scales exponentially with $L_{y}$. That said, for a given $L_{y}$, the entanglement entropy $S_{A}$ saturates at negative lower bound [73] $S_{min}\sim-L_{y}\log[a_{0}(b_{0}-1)^{-B}]$ because the system is ultimately gapped, such that the overlap $\eta$ at $k_{0}=\pi/L$ (nearest point to $k=0$) does not approach arbitrarily close to 1 with increasing $L$. For PBCs, the bulk gap also results in the saturation of $S_{A}$ at a positive value, as depicted by the starred grey trend in Fig. 2(e).
2. Gapless topological flat bands ($b_{0}=1$) with super volume-law negative entanglement. – Finally, we discuss the most intriguing case where $\det[H_{\text{y-OBC}}(k)]=0$ at $k=0$ [Eq. (11)], such that the topological band gap vanishes exactly [Fig. 3(a)]. Even though its band structure looks superficially similar to the $b_{0}>1$ case with exponentially small gap [Fig. 2], its entanglement entropy exhibits a surprising super volume-law dependence $S_{A}\sim-\frac{1}{2}B^{2}L_{y}^{2}\log L$, proportional not to the cylinder length $L_{y}$, but to the square of it. While the first power of $L_{y}$ can be attributed to the exponentially high topological band flatness as before, the additional second power of $L_{y}$ emerges from an uniquely new 2D phenomenon which we call non-Hermitian critical skin compression (nHCSC).
To understand the nHCSC, recall that all states are pushed to the cylinder’s edges due to the $y$-hopping asymmetry in our model Eq. (9). In particular, the left and right biorthogonal topological eigenstates exhibit exponential spatial-$y$ profiles $\bra{\psi_{edge}^{L}(k,y)}\sim r(k)^{-y}$ and |$\ket{\psi_{edge}^{R}(k,y)}\sim r(k)^{y}$ with $r(k)=\sqrt{|(1-\cos k)^{B}/a_{0}|}<1$, and are localized at edges $y=L_{y}$ and $y=1$ respectively. Since these are the two eigenstates that exhibit the highest overlap $\eta(k)$, they dominate the entanglement contribution in $\bar{P}$. Due to their criticality (gaplessness), $r(0)=0$ and the state is perfectly skin-localized for the $k=0$ topological mode in the $x$-direction along the cylinder edge.
Figure 3: Very robustly super volume-law negative entanglement entropy $S_{A}$ from gapless topological flat bands of our 2-component Hamiltonian (Eq. (9)) with parameters $b_{0}=1,t=0.5,a_{0}=2$. (a) Even for small $L_{y}=3,B=2$, a topological gapless flat band (bolded) with dispersion $E_{e}(k)\sim k^{BL_{y}}$ and overlap $\eta(k)\approx 1$ emerges around $k=0$ under $y$-OBCs. (b) For this gapless case, occupancy eigenvalues $p_{i}$ (Eq. 13) of $\bar{P}$ dramatically exceed the $[0,1]$ interval due to non-Hermitian critical skin compression (nHCSC), with $p_{1},p_{2},p_{3},...$ (red, green, blue…) exhibiting a hierarchy of power-law dependencies with $L$ (inset). (c) The negative entanglement scaling is accurately approximated by $S_{A}\approx-\frac{1}{2}(BL_{y})^{2}\log L$ (Eq. (14), black) across different $B,L_{y}$ combinations for sufficiently large cylinder circumference $L$. This is manifestly super volume-law behavior with respect to the cylinder length $L_{y}$. (d) The coefficient of $\log L$ in the numerical $S_{A}$, extracted through the gradient of the $dS_{A}/d(\log L)$ plots (shades of blue), agrees well with $-\frac{1}{2}(BL_{y})^{2}$ (Eq. (14) when $L\gtrapprox 10^{2}$. At smaller $L$, super volume-law dependence $\propto(BL_{y})^{2}$ still holds for smaller $BL_{y}$, albeit with a smaller coefficient.
Ordinarily, this perfect edge localization (or ”compression”) only leads to irreversible 1D non-Bloch dynamics [102] and singular generalized Brillouin zones  [103, 104, 105, 106, 107, 108, 109, 110]. However, in our 2D topological entanglement context, it also causes the occupied band projector
to diverge for matrix blocks $y<y^{\prime}$, with strongest divergence in $[P(k)]_{1,L_{y}}\sim r(k)^{-L_{y}}\sim(1-\cos k)^{-BL_{y}/2}$. Notably, the most strongly divergent contribution $\sim k^{-BL_{y}}$ from $(1-\cos k)^{-BL_{y}/2}$ does not dominate the total negative entanglement; of also substantial significance are the entire set of divergent terms $k^{-BL_{y}+2},k^{-BL_{y}+4},k^{-BL_{y}+6},...$ from the sub-leading terms in the expansion of $(1-\cos k)^{-BL_{y}/2}$, as well as other $[P(k)]_{y,y^{\prime}}$ [see [73] Sect. II.B]. Consequently, distinct from ordinary EP crossings [73, 65, 66], essentially the entire set of $\bar{P}$ eigenvalues $p_{i}$ diverges with $L$:
$p_{i}\sim\{L^{\frac{BL_{y}-1}{2}},L^{\frac{BL_{y}-1}{2}-1},L^{\frac{BL_{y}-1}{2}-2},...\}$ (13)
This is the main consequence of nHCSC, which hinges on both the edge compression of the eigenstates and its criticality (vanishing of $r(k)$). The hierarchy of these divergent occupancy eigenvalues is shown in Fig. 3(b): upon closer inspection, subdominant $p_{2},p_{3}\sim L^{\frac{BL_{y}-1}{2}-1},L^{\frac{BL_{y}-1}{2}-2}$ eigenvalues (green, blue) are observed in addition to the dominant $p_{1}\sim L^{\frac{BL_{y}-1}{2}}$. Summing over them, the total entanglement entropy scales like [73]
$S_{A}$ $=$ $-\sum_{p_{i}}p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i})$ (14)
$\approx$ $-\frac{1}{2}(BL_{y})^{2}\log L.$
This strongly negative entanglement $S_{A}$ is plotted in Fig. 3(c) for various $B,L_{y}$, and can be as low as $-70$ for reasonably large $L_{y}=4,B=2$. By examining the slope of $S_{A}$ with respect to the universal $\log L$ factor, it is numerically confirmed in Fig. 3(d) that the super volume-law quadratic coefficient $-(BL_{y})^{2}/2$ accurately holds for across a wide range of $BL_{y}$ as long as $L\gtrapprox 10^{2}$ (even though moderately large $L\sim\mathcal{O}(10)$ suffices when $BL_{y}$ is also of $\mathcal{O}(10)$).
Discussion.– We have established that negative free fermion entanglement fundamentally relies on substantial eigenstate overlap around the Fermi surface, which is less stringent than the previously suggested requirement of an exceptional crossing [65]. As such, macroscopically degenerate flatbands resulting from simultaneous topological and non-Hermitian skin localization can lead to strongly negative entanglement entropy, even in the presence of a small gap. Most strikingly, when this gap closes, the system furthermore exhibits super volume-law entanglement scaling $S_{A}\sim-\frac{1}{2}L_{y}^{2}\log L$, even though the entanglement cut is only of length $L_{y}$. This is attributed to a new mechanism known as non-Hermitian critical skin compression (nHCSC), where the criticality of highly degenerate NHSE-compressed topological modes gives rise to an extensive hierarchy of probability non-conserving $\bar{P}$ eigenstates that gives rise to even stronger negative $S_{A}$. This negative entanglement also applies to the Rényi entropy which can be physically measured as suggested in [73], placing topology as a potentially practical control knob for probability non-conserving negative entanglement.
Acknowledgments.– We thank Ruizhe Shen and Linhu Li for helpful discussions. This work is supported by the National Research Foundation, Singapore under its QEP2.0 programme (NRF2021-QEP2-02-P09) as well as the Ministry of Education, Singapore (MOE award number: MOE-T2EP50222-0003).
Supplemental information for ”Topologically protected negative entanglement”
This supplement contains the following material arranged by sections:
1. Detailed analysis of the different possible dispersions of the topological edge bands in a generalized 4-band exceptional topological model.
2. Analysis of the various scaling properties of the entanglement entropy $S_{A}$ in the topological flat band model, for both gapless and gapped cases.
3. How negativity of the entanglement entropy implies negative Rényi entropy, and constraints on the entanglement spectrum from symmetry.
4. A scheme for measuring negative entanglement through the second Rényi entropy.
I I. Detailed analysis of the 4-band exceptional topological model
I.1 A. General form of the Hamiltonian and the EP in its bulk bands
@@
$H(\mathbf{k})=\left(\sum_{j=x,y,z}\cos k_{j}-M\right)\tau_{z}\sigma_{0}+\lambda\sum_{j=x,y,z}\sin k_{j}\tau_{x}\sigma_{j}+[\sin\alpha\tau_{0}+\cos\alpha\tau_{z}](\vec{Z}\cdot\sigma)+i\delta\tau_{x}\sigma_{0}$ (S1)
@@
$\mathcal{H}(k,k_{z})$ $=$ $(\cos k+\cos k_{0}+\cos k_{z}-M)\tau_{z}\sigma_{0}+\lambda(\sin k\tau_{x}\sigma_{x}+\sin(-k_{0})\tau_{x}\sigma_{y}+\sin k_{z}\tau_{x}\sigma_{z})$ (S2)
$+$ $(\sin\alpha\tau_{0}+\cos\alpha\tau_{z})\vec{Z}\cdot\vec{\sigma}+i\delta\tau_{x}\sigma_{0},$
@@
$H(k,k_{y})$ $=$ $\big{(}\cos(k+k_{0})+\cos k_{0}+\cos k_{y}-M\big{)}\tau_{x}\sigma_{0}+\lambda\big{(}\sin(k+k_{0})\tau_{y}\sigma_{x}+\sin(-k_{0})\tau_{y}\sigma_{y}+\sin k_{y}\tau_{y}\sigma_{z}\big{)}$ (S3)
@@
I.1.1 Effect of EP dispersion on entanglement scaling
@@
In Figs. S1 and S2, we aim to demonstrate that the presence of an exceptional points (EP), whether topologically protected or not, does not necessarily imply the occurrence of a negative EE; the outcome also depends on the energy dispersion around the EP. In Fig. S1(a,b), we present a topologically non-trivial configuration, characterized by blue lines that represent the topological edge states for only OBCs and not PBCs. As contrasted with the linear dispersion around the EP discussed in the main text, this scenario exhibits a square-root dispersion, $E_{e}(k)\sim\sqrt{\delta k}$.
@@
Correspondingly, the overlap factor $\eta(k)=|\bra{\psi_{m}^{R}}\psi_{n}^{R}\rangle|^{2}/[\bra{\psi_{m}^{R}(k)}\psi_{m}^{R}(k)\rangle\bra{\psi_{n}^{R}(k)}\psi_{n}^{R}(k)]\rangle$ between the two topological states indexed by $m,n$ diminishes rapidly away from the EP $k=0$. It should be noted that in a tight-binding model, the momentum $k$ is discretized, with $k_{1}=\pi/L$ serving as the closest approximation to $k=0$. Given the rapid decay of $\eta(k)$ around $k=0$, it follows that $\eta(k_{1})$ does not approach 1. Consequently, as observed in Fig. S2(a), $\bar{P}$ exhibits only short-range hoppings, with its eigenvalues $p_{i}$ almost all located within the range $[0,1]$ (Fig. S2(b)) just like for an ordinary non-exceptional model (other than the fact that $p_{i}$ possess imaginary parts). Furthermore, the entanglement entropy remains positive and increases steadily with system size $L$, as shown in Fig. S2(c). Evidently, this scenario with $E_{e}(k)\sim\sqrt{\delta k}$ dispersion does not exhibit negative entanglement, suggesting that the dispersion around the EP crucially affects whether the entanglement entropy becomes negative. We will elaborate on how this dispersion can be tuned in our model in subsection I.B.
Figure S1: Parameter and boundary dependence of EPs in our 4-band model given by Eq. (7) of the main text. (a) For $\alpha=0$ and OBCs in the $y$ direction, an EP occurs at $k=0$. (b) With PBC in the $y$ direction, no EP is observed, implying that the EP arises due to boundary localization. (c) For $\alpha=\pi/2$ with OBCs in the $y$ direction, no EP exists either. Other parameters: $L_{y}=6,M=3,Z=0.44,\lambda=\delta=1$.
Figure S2: Absence of (a) divergent truncated projector $\bar{P}$ matrix elements and (b) eigenvalues outside of $[0,1]$ for an EP with square-root singularities. (c) Consequently, the real part of the entanglement entropy $S_{A}$ does not exhibit any negativity. Parameters and model are the same as in Fig.S1(a), with $x$-direction size $L=50$ and $L_{y}=6$.
@@
Figure S3: Slight negative dip in the entanglement entropy due to bulk EPs. Shown is the model in Eq. (7) of the main text, with the Fermi Surface fixed at $E_{F}=-1.1$ and parameters $L_{y}=3,M=3$ and $\delta=2$. (a) The truncated projector $\bar{P}$ matrix elements for $L=50$ with truncation at $l=L/2=25$. (b) The scaling of the maximum and minimum $p_{i}$ with system size $L$, which goes out of the $[0,1]$ interval only around $L=90$. (c) The entanglement entropy $S_{A}$ hence exhibits a slight negative dip at $L\approx 90$. However, this negative dip from the bulk EP is too weak to cause the entanglement to scale negatively as a whole.
I.1.2 Effect of bulk states on entanglement scaling due to topological EPs
@@
Returning to our model in Eq. (7) of the main text, we noted that EPs also exist in the bulk bands around Re$[E(k)]\approx\pm 1.1$. To investigate whether these bulk EPs could similarly induce a negative-valued entanglement entropy, we adjusted the Fermi energy to align with a bulk EP, setting $E_{F}$ to $1.1$. Observations from Fig.S3(a) indicate that the matrix elements of $\bar{P}$ remain minimal, with $p_{i}$ going beyond $[0,1]$ interval only around $L=90$ (see Figs.S3(b)). Consequently, the entanglement entropy exhibits a dip at this system size, albeit only a small dip. This occurs because the bulk EP is located at $k^{\prime}\approx 0.0106\pi$, rather than at the long-wavelength limit $k=0$. As $L$ increases, $k_{1}=\pi/L$ will pass through the EP $k^{\prime}$, inducing a dip in $S_{A}$. However, $S_{A}$ generally remains positive and continues to increase with increasing $L$, indicating that the negative scaling observed in Fig. 1(d) of the main text must be predominantly a consequence of the EP in the topological edge bands crossing at $k=0$, rather than of those within the bulk bands.
@@
I.2 B. Deriving two types of dispersion relations (linear and square-root) for topological edge bands

### v4

@@ II.1.1 Negative entanglement from exceptional topological crossing
In this section, we showcase a 4-band model featuring two topological edge modes that intersect at an exceptional crossing, rather than forming flat bands. We compare its entanglement scaling behavior with that of the flat-band model discussed in the Results, thereby providing a more comprehensive perspective on topologically protected negative entanglement. Unlike typical topological band crossings [73, 74] where the topological modes just have to be energetically degenerate, here we require them to also coalesce i.e. become parallel. A candidate model is given by the following 4-band Hamiltonian (see Methods, subsection A.2)
@@
In Fig. 3, we present three distinct scenarios corresponding to different parameter combinations, focusing particularly on the overlap $\eta(k)$ between the middle two eigenstates which straddle the Fermi energy $E_{F}=0$ (dashed line). Open boundary conditions (OBCs) are taken only along the $y$ direction, such that $k$ remains a good quantum number. In Fig. 3(a) with intersecting Hermitian topological modes (blue), $\eta(k)=0$ due to the exact orthonormality of Hermitian eigenstates. In Fig. 3(b) which is non-Hermitian ($\delta\neq 0$), $\eta(k)$ remains essentially zero due to the substantial line gap along Re$(E)=0$. However, in the non-Hermitian case with topological modes [Fig. 3(c)], the topological edge modes (red) cross and coalesce, forming an exceptional point, as reflected by the saturated squared overlap of $\eta(0)=1$. Only for this exceptional topological case do we see negatively entanglement entropy $\text{Re}(S_{A})$ [Fig. 3(d)]; for the previous two gapless and gapped cases of Figs.3(a,b), $\text{Re}(S_{A})$ respectively grows/saturates with $L$ as expected from usual conformal field theory [75, 76, 77].
@@
$\rm{Re}(S_{A})\sim-0.3399\log L\approx\left(\frac{1}{3}-\frac{2}{3}\right)\log L,$ (21)
@@
scaling in the exceptional topological case differs from the previously reported $S_{A}\sim-\frac{2}{3}\log L$ scaling for a linearly dispersive exceptional point [34, 35, 36]. This discrepancy is attributed to non-exceptional gapless crossing [gray in Fig. 3(c)], which contributes the usual $\frac{1}{3}\log L$ entanglement. As such, the negative entanglement from exceptional topological crossings can be easily overshadowed by other non-exceptional topological crossings, and is in this sense not necessarily robust  [^5] This is particularly so if the exceptional dispersion is square-root, as for some generalizations of our model in Methods, subsection A.1..
II.1.2 General form of the Hamiltonian and the EP in its bulk bands
Figure 4: Parameter and boundary dependence of EPs in our 4-band model given by Eq. (24). (a) For $\alpha=0$ and OBCs in the $y$ direction, an EP occurs at $k=0$. (b) With PBC in the $y$ direction, no EP is observed, implying that the EP arises due to boundary localization. (c) For $\alpha=\pi/2$ with OBCs in the $y$ direction, no EP exists either. Other parameters: $L_{y}=6,M=3,Z=0.44,\lambda=\delta=1$.
Figure 5: Absence of (a) divergent truncated projector $\bar{P}$ matrix elements and (b) eigenvalues outside of $[0,1]$ for an EP with square-root singularities. (c) Consequently, the real part of the entanglement entropy $S_{A}$ does not exhibit any negativity. Parameters and model are the same as in Fig.4(a), with $x$-direction size $L=50$ and $L_{y}=6$.
@@ II.1.2 General form of the Hamiltonian and the EP in its bulk bands
$H(\mathbf{k})$ $=$ $\left(\sum_{j=x,y,z}\cos k_{j}-M\right)\tau_{z}\sigma_{0}+\lambda\sum_{j=x,y,z}\sin k_{j}\tau_{x}\sigma_{j}$ (22)
$+$ $[\sin\alpha\tau_{0}+\cos\alpha\tau_{z}](\vec{Z}\cdot\sigma)+i\delta\tau_{x}\sigma_{0}$
@@
$\mathcal{H}(k,k_{z})$ $=$ $(\cos k+\cos k_{0}+\cos k_{z}-M)\tau_{z}\sigma_{0}+\lambda\big{(}\sin k\tau_{x}\sigma_{x}$ (23)
$+$ $\sin(-k_{0})\tau_{x}\sigma_{y}+\sin k_{z}\tau_{x}\sigma_{z}\big{)}+(\sin\alpha\tau_{0}+\cos\alpha\tau_{z})\vec{Z}\cdot\vec{\sigma}$
$+$ $i\delta\tau_{x}\sigma_{0},$
@@
$H(k,k_{y})$ $=$ $\big{(}\cos(k+k_{0})+\cos k_{0}+\cos k_{y}-M\big{)}\tau_{x}\sigma_{0}$ (24)
$+$ $\lambda\big{(}\sin(k+k_{0})\tau_{y}\sigma_{x}+\sin(-k_{0})\tau_{y}\sigma_{y}+\sin k_{y}\tau_{y}\sigma_{z}\big{)}$
@@
[…]
Effect of EP dispersion on entanglement scaling
@@
In Figs. 4 and 5, we aim to demonstrate that the presence of an exceptional points (EP), whether topologically protected or not, does not necessarily imply the occurrence of a negative EE; the outcome also depends on the energy dispersion around the EP. In Fig. 4(a,b), we present a topologically non-trivial configuration, characterized by blue lines that represent the topological edge states for only OBCs and not PBCs. As contrasted with the linear dispersion around the EP discussed in the main text, this scenario exhibits a square-root dispersion, $E_{e}(k)\sim\sqrt{\delta k}$.
@@
Correspondingly, the overlap factor $\eta(k)=|\bra{\psi_{m}^{R}}\psi_{n}^{R}\rangle|^{2}/[\bra{\psi_{m}^{R}(k)}\psi_{m}^{R}(k)\rangle\bra{\psi_{n}^{R}(k)}\psi_{n}^{R}(k)]\rangle$ between the two topological states indexed by $m,n$ diminishes rapidly away from the EP $k=0$. It should be noted that in a tight-binding model, the momentum $k$ is discretized, with $k_{1}=\pi/L$ serving as the closest approximation to $k=0$. Given the rapid decay of $\eta(k)$ around $k=0$, it follows that $\eta(k_{1})$ does not approach 1. Consequently, as observed in Fig. 5(a), $\bar{P}$ exhibits only short-range hoppings, with its eigenvalues $p_{i}$ almost all located within the range $[0,1]$ (Fig. 5(b)) just like for an ordinary non-exceptional model (other than the fact that $p_{i}$ possess imaginary parts). Furthermore, the entanglement entropy remains positive and increases steadily with system size $L$, as shown in Fig. 5(c). Evidently, this scenario with $E_{e}(k)\sim\sqrt{\delta k}$ dispersion does not exhibit negative entanglement, suggesting that the dispersion around the EP crucially affects whether the entanglement entropy becomes negative. We will elaborate on how this dispersion can be tuned in our model in Methods, subsection A.3.
@@
Figure 6: Slight negative dip in the entanglement entropy due to bulk EPs. Shown is the model in Eq. (7) of the main text, with the Fermi Surface fixed at $E_{F}=-1.1$ and parameters $L_{y}=3,M=3$ and $\delta=2$. (a) The truncated projector $\bar{P}$ matrix elements for $L=50$ with truncation at $l=L/2=25$. (b) The scaling of the maximum and minimum $p_{i}$ with system size $L$, which goes out of the $[0,1]$ interval only around $L=90$. (c) The real part of entanglement entropy $S_{A}$ hence exhibits a slight negative dip at $L\approx 90$. However, this negative dip from the bulk EP is too weak to cause the entanglement to scale negatively as a whole.
[…]
Effect of bulk states on entanglement scaling due to topological EPs
@@
Returning to our model in Eq. (7) of the main text, we noted that EPs also exist in the bulk bands around Re$[E(k)]\approx\pm 1.1$. To investigate whether these bulk EPs could similarly induce a negative-valued entanglement entropy, we adjusted the Fermi energy to align with a bulk EP, setting $E_{F}$ to $1.1$. Observations from Fig. 6(a) indicate that the matrix elements of $\bar{P}$ remain minimal, with $p_{i}$ going beyond $[0,1]$ interval only around $L=90$ (see Fig. 6(b)). Consequently, the entanglement entropy exhibits a dip at this system size, albeit only a small dip. This occurs because the bulk EP is located at $k^{\prime}\approx 0.0106\pi$, rather than at the long-wavelength limit $k=0$. As $L$ increases, $k_{1}=\pi/L$ will pass through the EP $k^{\prime}$, inducing a dip in $S_{A}$. However, $S_{A}$ generally remains positive and continues to increase with increasing $L$, indicating that the negative scaling observed in Fig. 1(d) of the main text must be predominantly a consequence of the EP in the topological edge bands crossing at $k=0$, rather than of those within the bulk bands.
@@
II.1.3 Deriving two types of dispersion relations (linear and square-root) for topological edge bands